OpenStudy (anonymous):

The sum of the first and second of thee consecutive odd integers is four more than four times the third integer. Find the third integer.

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

Something like that?

OpenStudy (anonymous):

Should be A + B = 4 + 4C

OpenStudy (anonymous):

Not A+B+4=4C

OpenStudy (anonymous):

thats what i'm double checking right now... but, we both did our work correctly! it just depends on how it's read

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

so who is right? @DemolisionWolf ?

OpenStudy (anonymous):

But solution will differ in values... For A+B=4+4C we have {-9,-7,-5} Whereas fot A+B+4=4C yields {-5,-3,-1}

OpenStudy (anonymous):

No.. sum of A+B IS Means A+B=

OpenStudy (anonymous):

Four more than what? Four more than 4C!!! So 4+ 4c

OpenStudy (anonymous):

Thus A+B=4+4C

OpenStudy (anonymous):

you are right aj.est1979, A+B=4+4C i'm gonna delete my post above